OpenStudy (df001):

The first term of a geometric sequence is 1 and the 4th term is 27. What is the 8th term?

OpenStudy (df001):

Help

OpenStudy (hockeychick23):

first term=1 4th term= 27 1*3= 3 2nd term 3*3=9 3rd term 9*3= 27 4th term 27*3= 81 5th term 81*3= 243 6th term 243*3= 729 7th term 729*3=2187 8th term